Citation and Reference List According to APA Manual

Erik Meijer

Department of Psychometrics and Research Methodology

Leiden University PO Box 9555 2300 RB Leiden The Netherlands E-Mail: Meijer@rulfsw.LeidenUniv.nl

June 7, 1999

1 Introduction

The American Psychological Association (APA) is very strict in the style submitted articles to its journals must obey. An important part of this style is the way citations should be typed. All requirements of the APA can be found in the Publication Manual of the American Psychological Association (3rd ed., 1983; revisions, 1984), Washington, DC: American Psychological Association, in the sequel just called the APA manual.

Several authors have written TEX and L^{A}TEX packages (.sty) and BibTEX style files (.bst) with suggestive names as apa, apalike, apalike2, newapa, theapa to try to capture some of the requirements of the APA manual. A severe problem with the APA citations is that, if the number of authors is between three and five, inclusive, the first citation must include all authors, whereas subsequent citations should refer to the first author “et al.”. This problem was solved by Young U. Ryu in his theapa package and bibstyle.

A lot of APA peculiarities, however, were not covered by these citation styles. Examples are:

• In the reference list, authors should be formatted with their initials after their surnames. Editors, however, should be formatted with their initials before their surnames if they are not in the author position.

• If the work referred to is a non-English-language source, the English translation of the title should be given in brackets after the title.

• Volume and edition of a book and the pages of an article in a book should be together between parentheses.

• The publisher of a book and its address should be given as “address: publisher”, and not as “publisher, address”.

• When the work used is a reprint or translation of an earlier work, this should be mentioned in the reference list and in text, the year must be given in the form “1923/1961”.

• There are several “tricky” things with brackets and parentheses, for example, with the type of an article (e.g., “Letter to the editor”, “Review of [some book]”, “Summary”, “Special issue”); magazines and newspapers are treated differently than journal articles; papers presented at meetings must be treated differently; translators of articles and books and chairs of symposia must be introduced; an article can have editors instead of authors (e.g., with a special issue); a PhD thesis can be abstracted in Dissertation Abstracts International; a (technical) report of some institution can be an (edited) collection of articles.

The apacite package and bibstyle are devised to get closer to the APA rules than the abovementioned other packages and styles.

The apacite.sty package consists for a large part of the citation part of theapa.sty. The main difference is that in apacite.sty, “no-parentheses” citation commands are provided. These can be used in tricky citations as “Merckens (1991; Bekker, Merckens, & Wansbeek, 1994) showed ....”,

which can be typed as “\citeauthor{Merc91} (\citeyearNP{Merc91};

\citeNP{BMW94}) showed ....”. Although this is not ideal, every (?) kind of in-text citation should now be possible.

By the way, it seems desirable that citation package writers use the same citation commands for the same purposes, and support them all. Because of the peculiarities of author-year type citations, a lot of extra commands have been introduced. If one would switch to one of the more standard type citation styles, sensible definitions of all these citation commands should be available. Concerning the different author-year citation styles, they should leave similar results with the same citation commands. For example, \citeA should always lead to citations of the form “author (year)”, and \cite should always lead to citations of the form “(author, year)” or vice versa.

Although the apacite.bst started out as a slightly adapted version of theapa.bst, the number of changes became so large and things became so unmanageable that I decided to write a completely new bibstyle, although using some small parts of the original. The apacite.bst file can therefore not anymore be seen as an adaptation of theapa.bst. This is the reason why the name is not theapa2.bst or similarly. Due to the inconsistent and complex rules of the APA manual, it has become a large file. In the process of writing,

the error You've exceeded BibTeX's wizard-defined function space

was very hard to overcome, but at last, it appears to be working. But, beware if you make changes (which are usually extensions)! You might run into the same problem again. (The solution is, of course, to recompile BibTEX with a larger function space, but this makes the style less portable.)

The final result, however, is not entirely capable of formatting every possible reference correctly, although the instances of wrong cases will be very rare. Many problems can be solved by “tricking” the style (e.g., by defining the number field as “2, Serial No. 186” if the issue number and serial number of a journal are both important), but this makes the .bib file less compatible with other styles. More problems, omissions, and things-to-be-done can be found in section 5.

In section 2, the in-text citation commands are described. Section 3 describes the format of the .bib file. Section 4 will discuss customization of the package.

2 Citations with apacite.sty

If the apacite package is loaded, the following citation commands can be used.

\cite<prefix>[postfix]{keys}
\fullcite<prefix>[postfix]{keys}
\shortcite<prefix>[postfix]{keys} These three commands write

\citation{keys} on the .aux file, and produce a citation of the following form: “(prefix Authors1, Year1; Authors2, Year2; ..., postfix)”, where the opening and closing parentheses and the commas and semicolons between author and year and between citations and between citations and postfix can be changed by the appropriate punctuation commands (see section 4). The command \fullcite uses the “full” author list, and the \shortcite command uses the abbreviated author list. The \cite command uses the “full” list the first time a work is referenced and the “short” list the second and subsequent times. The “full” and “short” lists are defined by the .bst file, and, in accordance with the APA manual, the “full” list consists of all authors if their number is five or less, and the first author suffixed with “et al.” (or whatever is defined by the command \BOthers) if their number is six or more. Similarly, the “short” list consists of all authors if their number is two or less and the first author with “et al.” if their number is three or more.

The prefix and postfix are optional. The “and” before the last author of a citation is defined by the command \BBAA.

\citeA[postfix]{keys}
\fullciteA[postfix]{keys}

\shortciteA[postfix]{keys} These commands are similar to the commands without the “A” suffix. They produce citations of the form “Authors1 (Year1), ..., AuthorsN (YearN, postfix)”. The parentheses are defined by the same commands as above, and the “and” before the last author of a citation is given by the command \BBAB.

\citeauthor{keys}
\fullciteauthor{keys}

\shortciteauthor{keys} These commands are similar to their *A counterparts, except that they produce citations of the form “Authors1, ..., AuthorsN” (the year is left out). The “and” before the last author is given by the \BBAB command.

\citeyear[postfix]{keys} Produces citations of the form “(Year1, ..., YearN, postfix)”.

\citeyearNP[postfix]{keys} A “no parentheses” version of \citeyear.

Produces citations of the form “Year1, ..., YearN, postfix”.

\citeNP<prefix>[postfix]{keys}
\fullciteNP<prefix>[postfix]{keys}
\shortciteNP<prefix>[postfix]{keys} “No parentheses” versions of

\cite, \fullcite, and \shortcite. They produce citations of the form “prefix Authors1, Year1; Authors2, Year2; ..., postfix”.

\nocite{keys} The entries are included in the bibliography without being cited in text. This is standard L^{A}TEX.

3 Making a .bib file for use with apacite.bst

To be able to obey the rules (and lots of exceptions) of the APA manual, lots of new entries (fields) and reference types had to be introduced. In section 3.1, the entries will be described, and in section 3.2, the reference types will be described.

3.1 Entries

The following entries can be used to describe a reference in the .bib file.

address^{†}^{1} The address (usually only the city) of the publisher, or school, or institution, or whatever published the item.

annote^{†} This is used in some annotated bibliography styles. It is not used by apacite, but no warning is given either. It is thus simply ignored.

author^{†} The author(s) of the work.

booktitle^{†} The title of the book or report in which an article was published.

chair The chair(s) of a symposium or meeting. Used for lectures. It is formatted the same way as editors.

chapter^{†} Not used. Although the APA manual defines an abbreviation for chapter, it is never used in the examples.

day The day of the month the referenced item was published (or whatever). Mainly used for magazine or newspaper articles or lectures.

edition^{†} The edition of the book or report. This must be of the form 1st, 2nd, 3rd, etc. Future work may provide routines to handle numbers only (1, 2, 3, etc.), so that language-specific texts are avoided.

editor^{†} The editor(s) of a book (or whatever).

englishtitle The English translation of an item with a non-English title.

firstkey The firstkey field, if not empty, is used as author for the first citation to an item (subsequent citations use the key field). This can be used if there is no author or editor field that can be used for citations, or in certain cases with corporate authors, where the citation in the text uses an abbreviation of the author’s name for second and subsequent citation, where the abbreviation is introduced in the first citation.

1The symbol^{†} means that the item is also described in the L^{A}TEXbook by Leslie Lamport and/or the L^{A}TEX Companion by Goossens, Mittelbach, and Samarin, although the specific meaning of the item may be (a little) different.

howpublished^{†} A description of how something was published, for example, “Unpublished manuscript” or “Paper presented at the Fourth International Conference on TEX Addiction”.

institution^{†} The institution or university (or whatever) that published a technical report.

journal^{†} The journal, magazine, newspaper, etc. in which an article or abstract of a thesis was published.

key^{†} See firstkey. The key field, if not empty, is used for for second and subsequent citations, or all citations if firstkey is missing. If author is empty and editor is empty or can not be used as author (e.g., in incollection), the key field is necessary to obtain a useful citation. This is so, because the APA manual requires an abbreviation of the title as in-text citation, and the abbreviation must be sensible. Therefore, in these cases, the user must provide one (note the quotes the APA manual requires).

month^{†} The month something was published. Mainly used for magazine or newspaper articles and lectures. Use the month macros jan, feb, etc.

note^{†} A note. This can be used for anything you like. See the APA manual for examples. Do not use it to denote the original publication of a reprint or translation, because the originalyear field must be used in those cases to obtain the correct citation (e.g., Freud, 1923/1961).

number^{†} The number of journal issue or report. This can be complex, for

example, “PRM 92-01”, or “5, Pt. 2”, or “10, Serial No. 134”.

Note that these latter examples imply that the generality (language and style independence) of the .bib file is somewhat lost.

organization^{†} Used for unpublished to denote the organization in which the unpublished work was produced.

originaladdress

originalbooktitle

originaledition originaleditor

originaljournal

originalnumber

originalpages

originalpublisher

originalvolume

originalyear These (original*-) fields have the same meaning as their counterparts without the “original” prefix, except that they refer to the book or journal in which the work was originally published. This can be used in referring to translations or reprints of articles in journals or books. If the original work is a book or other “standalone” work (except theses), only the originalyear field should be used. If the original work is a (PhD) thesis, which is also abstracted in Dissertation Abstracts Internation, or comparable, the originalyear field should be used to denote the year of the dissertation and the year field to denote the year of publication of the abstract.

pages^{†} The (inclusive) page numbers of the journal or book the article that is referred to was published in. Note that if a range of pages is given (which is usually the case), an en-dash should be used: “29--43”.

publisher^{†} The publisher of a book.

school^{†} The school or university for which a PhD thesis or master’s thesis was written.

series^{†} Not used.

symposium The name of the symposium or meeting at which a lecture was given.

text Used for items of type literal. This field contains the literal text to be used in the bibliography.

title^{†} The title of the work.

translator The translator of a book or article. This should be formatted in the

same way as author and editor. If the translator and editor fields

are identical, they are formatted as if there was only an editor, except that the editor receives the suffix “(Ed. and Trans.)” instead of just “(Ed.)”, or similarly if there is more than 1 editor, or if the commands

\BED, \BTRANS, \BEDS, \BTRANSS, and \BAnd are redefined.
type^{†} The type of phdthesis (e.g., Doctoral dissertation) or type
of article (e.g., Letter to the editor) or type of misc (e.g.,

Cassette recording), and similar types.

volume^{†} The volume or volumes of the book(s) referenced or the book in which the article was published, or the volume of the journal in which the article was published.

year^{†} The year in which the referenced item was published, or if it was not published, the year in which it was written. For accepted manuscripts or books that have not yet been published, “in press” can be used. The best way to do this is to use the “\BIP” command, which can be redefined by the user if the language or editorial style requires something else than “in press”.

3.2 Types of References

The work referenced to must be any of the following categories. In each category (or entry type) it is indicated which fields are relevant and which use they have. The firstkey and key fields can be used by each kind of entry, and the note field by each kind of entry except literal.

article^{†} A journal article or comparable. Obvious fields are author, year, title, journal, volume, pages. If the “article” referenced to is a special issue of a journal or something else that has editors instead of authors, the author field should be empty and the editor field should be used for the editors. If it is a non-English article, the English translation of the title should be given in the englishtitle field. If the journal paginates by issue instead of by volume, the issue number should be given in the number field. The type field can be used to denote the type of article, for example, “Letter to the editor”, or “Review of [some book or so]”. The translator field and original* fields can be used where appropriate.

magazine A magazine article. The fields that can be used are the same as with article, except that the month and day fields can also be used. The pages are formatted a little different (with “pp. 23--49” instead of just “23--49”).

newspaper This is equivalent to magazine. book^{†} An entire book. Obvious fields are author or editor, year,

title, volume, edition, address, publisher. The englishtitle,

translator, and original* fields can be used in the same way as

with article.

incollection^{†} An article in a (usually edited) book. Obvious fields are author,

year, title, editor, booktitle, volume, edition, pages, address,
publisher. The englishtitle, translator, and original* fields

can be used in the same way as with article.

techreport^{†} A report of some institution. Obvious fields are author, year, title, englishtitle, type (which is the type of report, e.g., “Technical report”, or “Cardiac arrest report”), number (the report number, e.g.,

“CA-94-005”), institution, address. The volume and edition fields

can also be used. If the report is an edited collection of articles, the editor field can be used instead of the author field.

intechreport An article in a (usually edited) report. Obvious fields are author, year, title, editor, booktitle (which is in this case the title of the entire report), pages, type (the type of report), number, institution,

address. The englishtitle, translator, and original* fields can

also be used.

lecture A paper presented at some meeting. Obvious fields are author, year, month, day (although APA manual mentions only month), title. The symposium field can be used for the name of the symposium and the chair field for the chair of the meeting. The howpublished field should be used to indicate the occasion at which the paper was presented if the symposium field is empty

(e.g., Paper presented at the First International Conference

on Bubblegum Marketing). It can also be used to give additional information about the symposium if the symposium field is not empty.

phdthesis^{†} A doctoral dissertation. For unpublished theses, the obvious fields

are author, year, title, englishtitle, school, address. The type

field can be used for the thesis type (e.g., “PhD thesis”). If it is empty, the default type (the command \BUPhD, which has the default value “Unpublished doctoral dissertation”) is used.

If the thesis is abstracted in Dissertation Abstracts International or similar, then the journal, volume, number, and pages fields can be used as with journal articles. The year field should denote the year of the ‘journal’. The originalyear field can be used to denote the year of the original thesis, if it is different from the year of the abstract. According to the APA manual, this should only be mentioned if the original thesis is used, and not the abstract. In this case, the school and type fields can be used as in the ‘unpublished thesis’ case, except that the type default is given by \BPhD, which has a default value of “Doctoral dissertation”.

mastersthesis^{†} This is equivalent to the “phdthesis” type, except that the default ‘unpublished’ and ‘published’ types are \BUMTh and \BMTh, with default values “Unpublished master’s thesis” and “Master’s thesis”.

unpublished^{†} For unpublished manuscripts and similarly ‘obscure’ material. Fields

that can be used are author or editor, year, month, day, title,
englishtitle, type and number. The howpublished field can be

used to indicate what kind of thing is referred to (e.g., “Unpublished manuscript”), and organization and address can be used to denote the organization (e.g., “Leiden University”) in which the document was produced and the address of the organization. These are formatted as “Leiden University, Leiden”.

misc^{†} For “things” that don’t fit the other categories, such as films, cassette recordings, and the like. Fields that can be used are author or editor,

title, year, month, day, type, number, howpublished, address, and

publisher. The address and publisher part is formatted as with books. The unpublished and misc types are very similar. The main difference

is that unpublished uses the organization field and misc uses the

publisher field, and they are formatted differently. Further, the

unpublished type requires author, editor, title, or type, whereas

with misc, howpublished takes the first position if these fields are all empty.

literal If the other categories do not format the item correctly, this category can be used. The text field is copied literally to the bibliography. The firstkey, key, year (and possibly originalyear) fields are necessary to get the correct in-text citations.

booklet^{†}

inbook^{†}

inproceedings^{†}

manual^{†} proceedings^{†} These categories are defined by all standard citation styles. They

are, however, not needed for the examples in the APA manual. For compatibility, however, they are included and defined as follows:

booklet = manual = proceedings = misc; inbook = inproceedings
= incollection.

4 Cite Punctuation and Labels

4.1 Punctuation

The following punctuation commands are provided and used for citations and reference list. They can be redefined by the user by \renewcommand.

\BBOP Open parenthesis, used for parentheses opening a citation, as in (Rao, 1973) or Rao (1973), and year field in reference list, as in

Rao, C. R. (1973). Default is “(”.

\BBAA Last “and” between authors in parentheses and reference section,

as in (Mooijaart \&{} Bentler, 1986) or Mooijaart, A., \&{}
Bentler, P. M. (1986). Default is “\&”.

\BBAB Last “and” between authors in text, as in Mooijaart and Bentler (1986). Default is “and”.

\BBAY Punctuation between author(s) and year in citation in parentheses, as in (Rao, 1973). Default is “, ”.

\BBC Punctuation between multiple cites, as in (Rao,1973;Mooijaart\&{}

Bentler, 1986). Default is “; ”.

\BBN Punctuation before note after citation, as in (Rao,1973,chapter 2). Default is “, ”.

\BBCP Closing parenthesis, used for parentheses closing a citation, as in (Rao, 1973) or Rao (1973), and year field in reference list, as in

Rao, C. R. (1973). Default is “)”.

\BBOQ Opening quote for article title in reference list, as in ``Random

polynomial factor analysis.'' Default is the empty string: no

quotes used.

\BBCQ Closing quote for article title in reference list. Default is the empty string. (Note that theapa.sty required the closing period to be part of the closing quote. This has been changed, because the period must be left out if title comments, such as type, English translation of non-English title, or translator follow the title.)

\BCBT Comma between authors in reference section when number of authors

     is two, as in “Mooijaart, A., \&{} Bentler, P. M. (1986).” The

comma after the “A.” is this one. Default is “,”.

\BCBL Comma before last author in citation and reference section,

     as in “(Gill, Murray, \&{} Wright, 1981)” or “Gill, P. E.,
     Murray, W., \&{} Wright, M. H. (1981).” Default is “,”.

If there are multiple citations with the same author and year, a letter should be added to the year to distinguish the references. For example, one may refer to two or more articles by J. Smith published in 1982. They should be referred to as “Smith, 1982a”, “Smith, 1982b”, and so forth. To accomplish this, the counter BibCnt is defined in apacite.sty, and the commands \BCnt and \BCntIP are defined, which convert the counter number to a letter. The command \BCntIP is used for multiple “in press” references with the same author (the “in press” should be defined by “\BIP”, see below), and the command \BCnt is used for other references. BibTEX calls these commands by \BCnt{number} or \BCntIP{number}, where number is simply 1, 2, …. Their definitions are

  \newcommand{\BCnt}[1]%
    {\setcounter{BibCnt}{#1}\alph{BibCnt}}

and

  \newcommand{\BCntIP}[1]%
    {\setcounter{BibCnt}{#1}{}-\alph{BibCnt}},

respectively. These commands can also be redefined by \renewcommand if desirable. For example,

  \renewcommand{\BCnt}[1]%
    {\setcounter{BibCnt}{#1}\emph{\alph{BibCnt}}}
  \renewcommand{\BCntIP}[1]%
    {\setcounter{BibCnt}{#1}{}-\emph{\alph{BibCnt}}}

would emphasize (italicize) the “a”, “b”, and so forth.

4.2 Labels

The following label commands are used in the reference section. They can be changed by the user by \renewcommand. These are (at least) language-specific, and therefore they are used in apacite.bst as commands, so as to improve the generality. \BOthers Used for “others” if number of authors is too large, as in

     “(Gill et~al., 1981)”. Default is “et~al.”.

\BIP “In press”, the string to be used as year for in-press-references. In the

.bib file, the year field should be “\BIP”. Default is “in press”. \BAnd Used as “and” for “editor and translator”, as in “In P. Watzlawick

     (Ed. and Trans.)”. Default is “and”.

\BED Editor in reference list, as in “In P. R. Krishnaiah (Ed.)”. Default

is “Ed.”. \BEDS Editors (more than one editor) in reference list, as in “In C. R. Rao

     \&{} J. K. Ghosh (Eds.)”. Default is “Eds.”.

\BTRANS Translator. Default is “Trans.”. \BTRANSS Translators. Default is “Trans.”. \BCHAIR Chair of a symposium, etc. Default is “Chair”. \BCHAIRS Chairs. Default is “Chairs”. \BVOL Volume, as in “Vol. 1”. Default is “Vol.”. \BVOLS Volumes, as in “Vols. 1--4”. Default is “Vols.”. \BNUM Number, as in “Tech. Rep. No. 12”. Default is “No.”. \BNUMS Numbers, as in “Nos. 3--5”. Default is “Nos.”. \BEd Edition, as in “2nd ed.”. Default is “ed.”. \BPG Page, default is “p.”. \BPGS Pages, default is “pp.”. \BTR The (default) technical report type name, which can be overridden by

the type entry. Default is “Tech.\ Rep.”

\BPhD The (default) PhD thesis type name, which can be overridden by the

     type entry. Default is “Doctoral dissertation”.

\BUPhD The (default) Unpublished PhD thesis type name, which can be overridden by the type entry. Default is “Unpublished doctoral

     dissertation”.

\BMTh The (default) master’s thesis type name, which can be overridden by the type entry. Default is “Master's thesis”.

\BUMTh The (default) unpublished master’s thesis type name, which can be overridden by the type entry. Default is “Unpublished master's

     thesis”.

\BOWP “Original work published”, default is “Original work published ”.

\BREPR “Reprinted from”, default is “Reprinted from ”.

The name of the bibliography itself (i.e., the heading) is defined by the commands \refname for articles and \bibname for books and reports. More specifically, \refname is used if \chapter is undefined, and \bibname is used otherwise. The commands \refname and \bibname can also be redefined by a \renewcommand. Default for both is “References”.

5 Problems, Things to be Done, etc.

The main problem is that the bibstyle needs to be “tricked” in many cases, which leads to correctly formatted references in APA style, but less compatibility of the .bib file with other styles (in case APA journals reject your article...). In the examples in the APA manual, frequently “double” references are required, for example, if a thesis is also abstracted in Dissertation Abstracts International, this should be reflected in the reference, and if it is obtained from microfilm or electronic document service, or whatever, this should also be mentioned.

Similar problems are encountered with reprints, translations, and works in languages other than English. With reprints and translations, the original work should also be mentioned (at least the year), with translations, the translator should be mentioned, and with non-English works, the English translation of the title should also be mentioned. Maybe the original-work problems can partially be solved by crossref-field-like cross reference, but cross referencing is absent from the current version.

Another class of problems is the numbering of journals and parts in journals. The usual way to refer to journal articles is to give the journal name, the volume, and the page number(s) of the article. If, however, a journal is paginated by issue, instead of by volume, the issue number must be given between parentheses directly following the volume. The issue number must be given in the number field. Sometimes, however, journals number their issues sequentially over volumes. In such cases, serial numbers can be given after the issue number per volume between the same parentheses: “Journal, 12 (2, Serial No. 46)”, where 12 is the volume number, 2 the issue number, and 46 the serial number. In this case the number field should be “2, Serial No. 46”, which is a kind of “tricking”. A similar problem is encountered when the article referred to is a monograph that forms a whole part of an issue. In such a case the number field should be, for example, “4, Pt. 2”. Maybe a part field, a serialno field, and so on solve some of these problems, but the amount of new fields necessary would cause the function space of BibTEX to overflow.

Many more of these problems exist, for example, with the ‘director’ and ‘producer’ of a film, or a ‘speaker’ of a cassette recording, or a ‘publisher’ and ‘distributor’ of all sorts of things. I haven’t even looked at legal cases (because I never refer to them and they don’t appear in the examples of the APA manual), but they will surely provide more problems.

Other known problems are:

entry size BibTEX predefines a global integer variable entry.max$ that sets the maximum length of an entry. In the implementation I have, this variable is set to 100. It is easy to imagine titles that are longer than 100 characters. Works with many authors or authors with long names may easily lead to an excessive author field. Moreover, I had

     to split the label field into full.label, label, and year.label to

be able to process the examples of the APA manual correctly (i.e., to keep below the maximum entry length). The most severe is this problem, however, with the sort.key$ field. Sorting can only be done on this field, and the complete author-year-title list can easily exceed 100 characters. Therefore, some tricks were implemented, one of which is that the entry list is sorted three times, where sort.key$ is, of course, defined differently in these cases.

authors with same last name References to J. Smith and P. Smith in the same article or book. The in-text citation should be (J. Smith, 1980; P. Smith, 1978), instead of (Smith, 1978, 1980). It is unclear to me if “A. Jones & J. Smith” and “H. Miller & P. Smith” should be referred to as “(Jones & J. Smith, 1980; Miller & P. Smith, 1978)”, or if simply “Jones & Smith” and “Miller & Smith” are allowed in this case.

I guess lots of other problems will come up when using this package/bibstyle combination, but at least I got it working on the APA manual examples. These examples and how they are processed by apacite can be found in the accompanying files testapa.tex and testapa.bib.